Supreme02:

Katarina drew a map to plot her running path as shown below. A running path is 0.5 + 2.2 + 0.9 + 3.2 inches long. If the total distance that she ran is 11.56 miles, how many miles does 1 inch on her map represent? Round to the nearest tenth if necessary. 0.4 miles 0.6 miles 1.7 miles 2.1 miles

jhonyy9:

so like a first step assume these numbers of 0.5 + 2.2 + 0.9 + 3.2 = ? after you get this result so just use the rule of 3 simple and write in this way 0.5 + 2.2 + 0.9 + 3.2 = ? inches mean 11,56 miles 1 in mean x miles and cross multiplie so calcule the x value x = (1*11,56)/(0.5 + 2.2 + 0.9 + 3.2 )= ? hope helped understandably easy @XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step assume these numbers of 0.5 + 2.2 + 0.9 + 3.2 = ? after you get this result so just use the rule of 3 simple and write in this way 0.5 + 2.2 + 0.9 + 3.2 = ? inches mean 11,56 miles 1 in mean x miles and cross multiplie so calcule the x value x = (1*11,56)/(0.5 + 2.2 + 0.9 + 3.2 )= ? hope helped understandably easy @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) I agree!!